### v1

@@ 3.2.2. $\alpha_{u}=\alpha_{v},\beta_{u}=\beta_{v}$
In the second case we fix some occurrence $v_{1}$ of $v$. There are at most $\frac{a+1}{2}$ overlays by 1’s of $v$ on $u$ such that $v_{short}^{-\alpha,\beta}$ coincides with $u_{short}^{-\alpha,\beta}$ (in particular, $v_{1}$ coincides with an equal subword). For each overlay there is a block $0^{x}$ in $v$ which is bigger than the corresponding block $0^{\alpha_{u}}$ in $u$. So, for each overlay we can add the block $0^{\alpha_{u}+1}$ to $v_{1}$ to the corresponding place, and for one of the overlays we take the block $0^{x}$ instead of the block $0^{\alpha_{u}+1}$. Let $v_{1}^{\prime}$ denote the obtained subword of $v$. The word $v_{1}^{\prime}$ is a subword of the word $v$ and is not a subword of $u$, and it contains $n_{1}$ occurrences of 1, at most $\frac{n_{short}-\alpha}{2}$ occurrences of 0 from small blocks, at most $\frac{a+1}{2}-1$ blocks $0^{\alpha+1}$ and one block $0^{x}$. Since $\alpha+1\leqslant x$, then
$|v_{1}^{\prime}|\leqslant n_{1}+\frac{(n_{short}-\alpha)}{2}+\frac{(a-1)(\alpha+1)}{2}+x=n_{1}+\frac{n_{short}}{2}+\frac{(a-2)(\alpha+1)+1}{2}+x\leqslant$
$\leqslant n_{1}+\frac{n_{short}}{2}+\frac{ax}{2}+\frac{1}{2}=n_{1}+\frac{n_{0}}{2}+\frac{1}{2}\leqslant\frac{3}{4}n+1.$
Since $|u_{1}^{\prime}|=|u_{1}|+x$ and $|v_{1}^{\prime}|\geqslant|u_{1}|+x$, then $|u_{1}^{\prime}|\leqslant\frac{3}{4}n+1$.
@@
(b) $u_{long}^{+\alpha,\beta}$ and $v_{long}^{+\alpha,\beta}$ are words of the third type
@@
In this case $a=b$ and $n_{short,u}=n_{short,v}=n_{short}$. Let $\beta_{u}\leqslant\alpha_{u},\alpha_{v},\beta_{v}$. Then $n_{1}=k(a+2)$. Consider the subword $u_{1}=(1^{p_{1}}0^{q_{1}}\cdots 1^{p_{i}}0^{q_{i}})1^{n_{1}-ks}$ of $u$. There are at most $\frac{k(a+2)}{ks}\leqslant\frac{a+2}{2}$ subwords of $u_{short}^{-\alpha,\beta}$ which are equal to $u_{1}$. Then either $v_{short}^{-\alpha,\beta}$ does not have a subword which is equal to $u_{1}$ or there is a subword $v_{1}=u_{1}$ of $v_{short}^{-\alpha,\beta}$. In the first case we add to $u_{1}$ a block $0^{x}$. Let $u_{1}^{\prime}$ denote the obtained subword of $u$. The word $u_{1}^{\prime}$ is a subword of the word $u$ and is not a subword of $v$. We will later prove that $|u_{1}^{\prime}|\leqslant\frac{3}{4}n+4$.
@@
In the second case we fix some occurrence $v_{1}$ of $v$, where all 0’s of this occurrence are taken from small blocks, except for $0^{\alpha_{v}}$ and $0^{\beta_{v}}$. There are at most $\frac{a+2}{2}$ overlays by 1’s of $v$ on $u$ such that $v_{short}^{-\alpha,\beta}$ coincides with $u_{short}^{-\alpha,\beta}$ (in particular, $v_{1}$ coincide with equal to $v_{1}$ subword). For each overlay there is a block $0^{x}$ in $v$ which is bigger than the corresponding block $0^{\alpha_{u}}$ or $0^{\beta_{u}}$ in $u$. So, for each overlay we can add the block $0^{\alpha_{u}+1}$ or $0^{\beta_{u}+1}$ to $v_{1}$ to the corresponding place. Notice that there is at least one overlay for which we can add the block $0^{\beta_{u}+1}$ (which is not the case for $0^{\alpha_{u}+1}$). And in the resulting subword we take one block $0^{x}$ instead of one of the blocks $0^{\alpha_{u}+1}$ (or instead of one of the blocks $0^{\beta_{u}+1}$ if we did not add the block $0^{\alpha_{u}+1}$). Let $v_{1}^{\prime}$ denote the obtained subword of $v$. The word $v_{1}^{\prime}$ is a subword of the word $v$ and is not a subword of $u$ and it contains $n_{1}$ occurrences of 1, at most $\frac{n_{short}-\alpha_{v}-\beta_{v}}{2}=\frac{n_{short}-\alpha_{u}-\beta_{u}}{2}$ occurrences of 0 from small blocks, one block $0^{\beta+1}$, one block $0^{x}$ and at most $\frac{a+2}{2}-2$ blocks $0^{\alpha_{u}+1}$ or $10^{\beta_{u}}+$ 1. Since $\beta_{u}+1\leqslant\alpha_{u}+1\leqslant x$, then
$|v_{1}^{\prime}|\leqslant n_{1}+\frac{(n_{short}-\alpha_{u}-\beta_{u})}{2}+x+\beta+1+\frac{(a-2)(\alpha+1)}{2}=$
$=n_{1}+\frac{n_{short}}{2}+\frac{2x+(\beta+1)+(a-3)(\alpha+1)+2}{2}\leqslant$
$\leqslant n_{1}+\frac{n_{short}}{2}+\frac{ax}{2}+1=n_{1}+\frac{n_{0}}{2}+1\leqslant\frac{3}{4}n+1.$
Since $|u_{1}^{\prime}|=|u_{1}|+x$ and $|v_{1}^{\prime}|\geqslant|u_{1}|+x$, then $|u_{1}^{\prime}|\leqslant\frac{3}{4}n+1$.
3.2. $s=1$.
@@
Let $w$ be a subword of only one of the words $u$ and $v$. Suppose that $w$ has $n_{1}$ occurrences of 1, at least one block $0^{x}$ and at most $\frac{2}{3}ax+\frac{4}{5}(\alpha_{u}+\beta_{u})+\frac{23}{10}$ occurrences of 0 from $t$ blocks of 0’s. If $t\geqslant\frac{2}{3}a$, then there is a subword of only one of the words $u$ and $v$ of length at most $\frac{3}{4}n+4$.
Proof 4.8.
Since $\alpha_{u}+\beta_{u}=\alpha_{v}+\beta_{v}$, then without of loss generality we can assume that $w$ is a subword of the word $u$ and is not a subword of $v$. Notice that there are at least $\frac{1}{3}ax+\frac{1}{5}(\alpha_{u}+\beta_{u})+(l-(a+2))-\frac{23}{10}$ occurrences of 0 which do not belong to $w$. If $|w|\leqslant\frac{3}{4}n+4$, then $w$ itself is a desired subword. Otherwise
@@
Multiplying it by 3, we get an equivalent inequality
@@
Consider the word $w^{\prime}=(01)^{l}$ and fix some its occurrence in $u$. We add to $w^{\prime}$ the same blocks of 0’s as in $w$, and we let $w^{\prime\prime}$ denote the obtained subword of $u$. Since $u_{short}^{-\alpha,\beta}$, $v_{short}^{-\alpha,\beta}$, $u_{long}^{+\alpha,\beta}$, $v_{long}^{+\alpha,\beta}$ are periodic with period $k$, then $w^{\prime\prime}$ is a subword of the word $u$ and is not a subword of $v$. Moreover,
@@
$=(ax+\frac{3}{5}(\alpha_{u}+\beta_{u})+3l-3a)-(l-a)-\left(\frac{1}{3}ax-\frac{4}{3}a-\frac{1}{5}(\alpha_{u}+\beta_{u})+t\right)+\frac{23}{10}<\frac{3}{4}n+4,$
@@
where the latter inequalities follow from the inequalities $t\geqslant\frac{2}{3}a$ and $\alpha_{u},\beta_{u}\leqslant x$. The claim is proved.
If $u_{long}^{+\alpha,\beta}=v_{long}^{+\alpha,\beta}$, then $u=v$ since $u_{short}^{-\alpha,\beta}\cong v_{short}^{-\alpha,\beta}$ and $u_{short}^{-\alpha,\beta}$ is periodic and length of period is $k$. A contradiction. So, $u_{long}^{+\alpha,\beta}\neq v_{long}^{+\alpha,\beta}$. Let
$u_{long}^{+\alpha,\beta}=0^{\alpha_{u}}1^{k}(0^{x}1^{k})^{r_{1}-1}0^{\beta_{u}}1^{k}(0^{x}1^{k})^{s_{1}-1},$
$v_{long}^{+\alpha,\beta}=0^{\alpha_{v}}1^{k}(0^{x}1^{k})^{r_{2}-1}0^{\beta_{v}}1^{k}(0^{x}1^{k})^{s_{2}-1}.$
@@
3.2.1. $\beta_{u}<\alpha_{u}$ and $\beta_{u}+1\leqslant\frac{2}{3}x$.
Consider a subword $v_{1}$ of $v$ which contains $n_{1}$ occurrences of 1 and $\beta_{u}+1$ occurrences of 0 from each of the blocks $0^{x}$ and from the block $0^{\alpha_{v}}$. Fix some occurrence of $v_{1}$ in $v$. There is at most one overlay by 1’s of $v$ on $u$ such that small blocks overlay on small blocks and $v_{1}$ coincides with an equal word. If such overlay is exists, then there is a block $0^{x}$ in $v$ which overlay on block $0^{\alpha_{u}}$ or $0^{\beta_{u}}$. We add this block $0^{x}$ to $v_{1}$. Otherwise we add any block $0^{x}$ to $v_{1}$. Let $v_{1}^{\prime}$ denote the obtained subword of $v$. Since $\beta_{u}+1\leqslant\alpha_{u}$ and $\beta_{u}+1\leqslant\frac{2}{3}x$, then the number of occurrences of 0 in $v_{1}^{\prime}$ is
@@
Note that $v_{1}^{\prime}$ has 0’s from at least $a$ blocks of 0’s. So, all conditions from the claim above hold. Hence there is a subword of only one of the words $u$ and $v$ of length at most $\frac{3}{4}n+4$.
In the following cases we have either $\beta_{u}=\alpha_{u}$ or $\beta_{u}+1>\frac{2}{3}x$. In both cases we have $\beta_{u}+1>\frac{2}{3}\alpha_{u}-1$. Besides that, since $\alpha_{u}\geqslant\alpha_{v}\geqslant\beta_{v}\geqslant\beta_{u}$, then $\beta_{v}+1>\frac{2}{3}\alpha_{v}$.
3.2.2. $\alpha_{u}>\alpha_{v}$.
Since $\alpha_{u}+\beta_{u}=\alpha_{v}+\beta_{v}$, then $\alpha_{u}>\alpha_{v}\geqslant\beta_{v}>\beta_{u}$. Let $u_{1}$ be the subword of $u$ containing $n_{1}$ occurrences of 1. We index blocks of 0’s in $u_{long}^{+\alpha,\beta}$ clockwise, and we split all blocks into groups such that each group contains $t$ blocks with indices $i,i+r_{2},i+2r_{2},\ldots,i+(t-1)r_{2}$ for some $i$. We add $\alpha_{v}+1$ zeros to $u_{1}$ from every second block from each group (either from blocks with indices $i,i+2r_{2},i+4r_{2},\ldots$ or from blocks with indices $i+r_{2},i+3r_{2},i+5r_{2},\ldots$). It is possible since there is only one block of 0’s in $u$ which is smaller than $0^{\alpha_{v}+1}$ (this is a block $0^{\beta_{u}}$). We added $\lceil\frac{t}{2}\rceil\leqslant\frac{2}{3}$ blocks of 0’s from each group, that is, at most $\frac{2(a+2)}{3}$ blocks of $0^{\alpha_{v}+1}$. We also replaced one block $0^{\alpha_{v}+1}$ with $0^{x}$. Without loss of generality, we can assume that we added 0’s exactly from $\frac{2(a+2)}{3}$ blocks. Let $u_{1}^{\prime}$ denote the obtained subword of $u$. The word $u_{1}^{\prime}$ is a subword of the word $u$ and is not a subword of $v$. Moreover, since $\alpha_{v}+1\leqslant x$ and $\beta_{v}+1\geqslant\frac{2}{3}\alpha_{v}$, then $u_{1}^{\prime}$ has at most
$\left(\frac{2(a+2)}{3}-1\right)(\alpha_{v}+1)+x\leqslant\frac{2}{3}ax+\frac{4}{3}(\alpha_{v}+1)\leqslant\frac{2}{3}ax+\frac{4}{5}(\alpha_{v}+\beta_{v}+1)+\frac{4}{3}$
occurrences of 0. The last inequality holds since
@@
Since $\frac{4}{5}+\frac{4}{3}<2,3$, then all conditions from the above claim hold. Hence, there is a subword of only one of the words $u$ and $v$ of length at most $\frac{3}{4}n+4$.
3.2.3. $\alpha_{u}=\alpha_{v}=\beta_{u}=\beta_{v}$.
@@
Since $u\neq v$, then $r_{1}\neq r_{2},s_{2}$ and $s_{1}\neq r_{2},s_{2}$. We index blocks of 0’s in $u_{long}^{+\alpha,\beta}$ and split all blocks into groups as in case 3.2.2. in a way that each group contains $t$ blocks with indices $i,i+r_{2},i+2r_{2},\ldots,i+(t-1)r_{2}$ for some $i$. If the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ are in the different groups, then we can add $\alpha_{v}+1$ zeros to $u_{1}$ from every second block from each group and proceed with the proof as in case 3.2.2.. Assume that blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ are in the same group, and the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ have indices $i$ and $i+hr_{2}$, respectively.
Since $\{r_{1},s_{1}\}\neq\{r_{2},s_{2}\}$, then $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ are not adjacent blocks in their group. That is $h\neq 1$ and $h\neq t-1$. In particular, $t\geqslant 4$. From each group which does not contain the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$, we can add $0^{x}$ to $u_{1}$ from every second block (from blocks with indices $i,i+2r_{2},i+4r_{2},\ldots$). If $h$ is even, then for the group containing the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ we can add $0^{x}$ to $u_{1}$ from blocks with indices $i+r_{2},i+3r_{2},i+5r_{2},\ldots$. If $h$ is odd, then for the group containing the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ we can add $0^{x}$ to $u_{1}$ from blocks with indices $i+r_{2},i+3r_{2},\ldots,i+(h-2)r_{2},i+(h-1)r_{2},i+(h+1)r_{2},i+(h+3)r_{2},i+(h+5)r_{2},\ldots$. It is not difficult to verify that for any $t\geqslant 4$ we add at most $\frac{2}{3}(a+2)$ blocks $0^{x}$ to $u_{1}$. We let $u_{1}^{\prime}$ denote the obtained subword of $u$. The proof in this case can be completed similarly to the case 3.2.2..
3.2.4. $\alpha_{u}=\alpha_{v}>\beta_{u}=\beta_{v}$.

### v2

@@ 4.2 Lemmas constituting the proof
In the second case we fix some occurrence of $v_{1}$ in $v$. There are at most $\frac{a+1}{2}$ different 1-overlays of $v$ on $u$ such that $v_{\operatorname{short}}^{-\alpha,\beta}$ coincides with $u_{\operatorname{short}}^{-\alpha,\beta}$ (in particular, $v_{1}$ coincides with an occurrence of an equal subword). For each 1-overlay there is a block $0^{x}$ in $v$ which is bigger than the corresponding block $0^{\alpha_{u}}$ in $u$. So, for each 1-overlay we can add the block $0^{\alpha_{u}+1}$ to $v_{1}$ to the corresponding place, and for one of the overlays we take the block $0^{x}$ instead of the block $0^{\alpha_{u}+1}$. Let $v_{1}^{\prime}$ denote the obtained subword of $v$. The word $v_{1}^{\prime}$ is a subword of the word $v$ and is not a subword of $u$, and it contains $n_{1}$ occurrences of 1, at most $\frac{n_{\operatorname{short}}-\alpha}{2}$ occurrences of 0 from small blocks, at most $\frac{a+1}{2}-1$ blocks $0^{\alpha+1}$ and one block $0^{x}$. Since $\alpha+1\leqslant x$, then
$|v_{1}^{\prime}|\leqslant n_{1}+\frac{(n_{\operatorname{short}}-\alpha)}{2}+\frac{(a-1)(\alpha+1)}{2}+x=n_{1}+\frac{n_{\operatorname{short}}}{2}+\frac{(a-2)(\alpha+1)+1}{2}+x\leqslant$
$\leqslant n_{1}+\frac{n_{\operatorname{short}}}{2}+\frac{ax}{2}+\frac{1}{2}=n_{1}+\frac{n_{0}}{2}+\frac{1}{2}\leqslant\frac{3}{4}n+1.$
Since $|u_{1}^{\prime}|=|u_{1}|+x$ and $|v_{1}^{\prime}|\geqslant|u_{1}|+x$, we have $|u_{1}^{\prime}|\leqslant\frac{3}{4}n+1$.
@@
Case 3.1.2. $u_{\operatorname{long}}^{+\alpha,\beta}$ and $v_{\operatorname{long}}^{+\alpha,\beta}$ are words of the third type.
@@
In this case $a=b$ and $n_{\operatorname{short},u}=n_{\operatorname{short},v}=n_{\operatorname{short}}$. Let $\beta_{u}\leqslant\alpha_{u},\alpha_{v},\beta_{v}$. Then $n_{1}=k(a+2)$. Consider the subword $u_{1}=(1^{p_{1}}0^{q_{1}}\cdots 1^{p_{i}}0^{q_{i}})1^{n_{1}-ks}$ of $u$. There are at most $\frac{k(a+2)}{ks}=\frac{a+2}{2}$ ways to choose an occurrence of a subword of $u_{\operatorname{short}}^{-\alpha,\beta}$ which is equal to $u_{1}$. Then either $v_{\operatorname{short}}^{-\alpha,\beta}$ does not have a subword which is equal to $u_{1}$ or there is a subword $v_{1}=u_{1}$ of $v_{\operatorname{short}}^{-\alpha,\beta}$. In the first case we add to $u_{1}$ a block $0^{x}$; and we let $u_{1}^{\prime}$ denote the obtained subword of $u$. The word $u_{1}^{\prime}$ is a subword of the word $u$ and is not a subword of $v$. We later prove that $|u_{1}^{\prime}|\leqslant\frac{3}{4}n+4$.
@@
In the second case we fix some occurrence $v_{1}$ of $v$, where all 0’s of this occurrence are taken from small blocks, except for $0^{\alpha_{v}}$ and $0^{\beta_{v}}$ (we do not take any 0’s from these blocks). There are at most $\frac{a+2}{2}$ different 1-overlays of $v$ on $u$ such that $v_{\operatorname{short}}^{-\alpha,\beta}$ coincides with $u_{\operatorname{short}}^{-\alpha,\beta}$ (in particular, $v_{1}$ coincides with an occurrence of a subword equal to $u_{1}$). For each 1-overlay there is a block $0^{x}$ in $v$ which is bigger than the corresponding block $0^{\alpha_{u}}$ or $0^{\beta_{u}}$ in $u$. So, for each 1-overlay we can add the block $0^{\alpha_{u}+1}$ or $0^{\beta_{u}+1}$ to $v_{1}$ to the corresponding place. Notice that there is at least one 1-overlay for which we can add the block $0^{\beta_{u}+1}$ (which is not the case for $0^{\alpha_{u}+1}$). Besides that, in the resulting subword we take one block $0^{x}$ instead of one of the blocks $0^{\alpha_{u}+1}$ (or instead of one of the blocks $0^{\beta_{u}+1}$ if we did not add the block $0^{\alpha_{u}+1}$). Let $v_{1}^{\prime}$ denote the obtained subword of $v$. The word $v_{1}^{\prime}$ is a subword of the word $v$ and is not a subword of $u$ and it contains $n_{1}$ occurrences of 1, at most $\frac{n_{\operatorname{short}}-\alpha_{v}-\beta_{v}}{2}=\frac{n_{\operatorname{short}}-\alpha_{u}-\beta_{u}}{2}$ occurrences of 0 from small blocks, one block $0^{\beta+1}$, one block $0^{x}$ and at most $\frac{a+2}{2}-2$ blocks $0^{\alpha_{u}+1}$ or $0^{\beta_{u}}+1$. Since $\beta_{u}+1\leqslant\alpha_{u}+1\leqslant x$, we have
$|v_{1}^{\prime}|\leqslant n_{1}+\frac{(n_{\operatorname{short}}-\alpha_{u}-\beta_{u})}{2}+x+\beta+1+\frac{(a-2)(\alpha+1)}{2}=$
$n_{1}+\frac{n_{\operatorname{short}}}{2}+\frac{2x+(\beta+1)+(a-3)(\alpha+1)+2}{2}\leqslant$
$\leqslant n_{1}+\frac{n_{\operatorname{short}}}{2}+\frac{ax}{2}+1=n_{1}+\frac{n_{0}}{2}+1\leqslant\frac{3}{4}n+1.$
Since $|u_{1}^{\prime}|=|u_{1}|+x$ and $|v_{1}^{\prime}|\geqslant|u_{1}|+x$, we have $|u_{1}^{\prime}|\leqslant\frac{3}{4}n+1$.
Case 3.2. $s=1$.
@@
Under the conditions of Case 3.2, let $w$ be a distinguishing subword for the words $u$ and $v$. Suppose that $w$ has $n_{1}$ occurrences of 1, at least one block $0^{x}$ and at most $\frac{2}{3}ax+\frac{4}{5}(\alpha_{u}+\beta_{u})+\frac{23}{10}$ occurrences of 0 from $t$ blocks of 0’s. If $t\geqslant\frac{2}{3}a$, then there is a distinguishing subword for the words $u$ and $v$ of length at most $\frac{3}{4}n+4$.
Proof 4.7.
Since $\alpha_{u}+\beta_{u}=\alpha_{v}+\beta_{v}$, without loss of generality we can assume that $w$ is a subword of the word $u$ and is not a subword of $v$. Notice that there are at least $\frac{1}{3}ax+\frac{1}{5}(\alpha_{u}+\beta_{u})+(l-(a+2))-\frac{23}{10}$ occurrences of 0 which do not belong to $w$. If $|w|\leqslant\frac{3}{4}n+4$, then $w$ itself is a desired subword. Otherwise
@@
Multiplying the inequality by 3, we get an equivalent inequality
@@
Consider the word $w^{\prime}=(01)^{l}$ and fix some occurrence of it in $u$. We add to $w^{\prime}$ the same blocks of 0’s as in $w$, and we let $w^{\prime\prime}$ denote the obtained subword of $u$. Since $u_{\operatorname{short}}^{-\alpha,\beta}$, $v_{\operatorname{short}}^{-\alpha,\beta}$, $u_{\operatorname{long}}^{+\alpha,\beta}$, $v_{\operatorname{long}}^{+\alpha,\beta}$ are periodic with period $k$, $w^{\prime\prime}$ is a subword of the word $u$ and is not a subword of $v$. Moreover,
@@
$(ax+\frac{3}{5}(\alpha_{u}+\beta_{u})+3l-3a)-(l-a)-\left(\frac{1}{3}ax-\frac{4}{3}a-\frac{1}{5}(\alpha_{u}+\beta_{u})+t\right)+\frac{23}{10}<\frac{3}{4}n+4,$
@@
[…]
If $u_{\operatorname{long}}^{+\alpha,\beta}=v_{\operatorname{long}}^{+\alpha,\beta}$, then $u=v$ since $u_{\operatorname{short}}^{-\alpha,\beta}\cong v_{\operatorname{short}}^{-\alpha,\beta}$ and $u_{\operatorname{short}}^{-\alpha,\beta}$ is periodic and the length of period is $k$. A contradiction. So, $u_{\operatorname{long}}^{+\alpha,\beta}\neq v_{\operatorname{long}}^{+\alpha,\beta}$. Let
$u_{\operatorname{long}}^{+\alpha,\beta}=0^{\alpha_{u}}1^{k}(0^{x}1^{k})^{r_{1}-1}0^{\beta_{u}}1^{k}(0^{x}1^{k})^{s_{1}-1},$
$v_{\operatorname{long}}^{+\alpha,\beta}=0^{\alpha_{v}}1^{k}(0^{x}1^{k})^{r_{2}-1}0^{\beta_{v}}1^{k}(0^{x}1^{k})^{s_{2}-1}.$
@@
Case 3.2.1. $\beta_{u}<\alpha_{u}$ and $\beta_{u}+1\leqslant\frac{2}{3}x$.
Consider the subword $v_{1}$ of $v$ which contains $n_{1}$ occurrences of 1 and $\beta_{u}+1$ occurrences of 0 from each of the blocks $0^{x}$ and from the block $0^{\alpha_{v}}$. Fix some occurrence of $v_{1}$ in $v$. There is at most one 1-overlay of $v$ on $u$ such that small blocks overlay on small blocks and $v_{1}$ coincides with an occurrence of an equal subword. If such a 1-overlay exists, then there is a block $0^{x}$ in $v$ which overlays on a block $0^{\alpha_{u}}$ or $0^{\beta_{u}}$. We add this block $0^{x}$ to $v_{1}$. Otherwise we add any block $0^{x}$ to $v_{1}$. Let $v_{1}^{\prime}$ denote the obtained subword of $v$. Since $\beta_{u}+1\leqslant\alpha_{u}$ and $\beta_{u}+1\leqslant\frac{2}{3}x$, the number of occurrences of 0 in $v_{1}^{\prime}$ is
@@
Note that $v_{1}^{\prime}$ has 0’s from at least $a$ blocks of 0’s. So, all conditions from the claim above hold. Hence there is a distinguishing subword for the words $u$ and $v$ of length at most $\frac{3}{4}n+4$.
In the following cases we have either $\beta_{u}=\alpha_{u}$ or $\beta_{u}+1>\frac{2}{3}x$. In both cases we have $\beta_{u}+1>\frac{2}{3}\alpha_{u}-1$. Besides that, since $\alpha_{u}\geqslant\alpha_{v}\geqslant\beta_{v}\geqslant\beta_{u}$, we have $\beta_{v}+1>\frac{2}{3}\alpha_{v}$.
Case 3.2.2. $\alpha_{u}>\alpha_{v}$.
Since $\alpha_{u}+\beta_{u}=\alpha_{v}+\beta_{v}$, we have $\alpha_{u}>\alpha_{v}\geqslant\beta_{v}>\beta_{u}$. We start building a distinguishing subword by taking $u_{1}$ to be the subword of $u$ containing $n_{1}$ occurrences of 1. We index blocks of 0’s in $u_{\operatorname{long}}^{+\alpha,\beta}$ clockwise, and we split all blocks into groups such that each group contains $t$ blocks with indices $i,i+r_{2},i+2r_{2},\ldots,i+(t-1)r_{2}$ for some $i$. We add $\alpha_{v}+1$ zeros to $u_{1}$ from every other block from each group (either from blocks with indices $i,i+2r_{2},i+4r_{2},\ldots$ or from blocks with indices $i+r_{2},i+3r_{2},i+5r_{2},\ldots$). It is possible since there is only one block of 0’s in $u$ which is smaller than $0^{\alpha_{v}+1}$ (this is a block $0^{\beta_{u}}$). We added $\lceil\frac{t}{2}\rceil\leqslant\frac{2}{3}$ blocks of 0’s from each group, that is, at most $\frac{2(a+2)}{3}$ blocks of $0^{\alpha_{v}+1}$. We also replaced one block $0^{\alpha_{v}+1}$ with $0^{x}$. Without loss of generality, we can assume that we added 0’s exactly from $\frac{2(a+2)}{3}$ blocks. Let $u_{1}^{\prime}$ denote the obtained subword of $u$. The word $u_{1}^{\prime}$ is a subword of the word $u$ and is not a subword of $v$. Moreover, since $\alpha_{v}+1\leqslant x$ and $\beta_{v}+1\geqslant\frac{2}{3}\alpha_{v}$, the number of 0’s in $u_{1}^{\prime}$ has at most
$\left(\frac{2(a+2)}{3}-1\right)(\alpha_{v}+1)+x\leqslant\frac{2}{3}ax+\frac{4}{3}(\alpha_{v}+1)\leqslant\frac{2}{3}ax+\frac{4}{5}(\alpha_{v}+\beta_{v}+1)+\frac{4}{3}.$
The last inequality holds since
@@
Since $\frac{4}{5}+\frac{4}{3}<2.3$, all conditions from the above claim hold. Hence, there is a distinguishing subword for the words $u$ and $v$ of length at most $\frac{3}{4}n+4$.
Case 3.2.3. $\alpha_{u}=\alpha_{v}=\beta_{u}=\beta_{v}$.
@@
Since $u\neq v$, we have $r_{1}\neq r_{2},s_{2}$ and $s_{1}\neq r_{2},s_{2}$. We index blocks of 0’s in $u_{\operatorname{long}}^{+\alpha,\beta}$ and split all blocks into groups as in Case 3.2.2 in a way that each group contains $t$ blocks with indices $i,i+r_{2},i+2r_{2},\ldots,i+(t-1)r_{2}$ for some $i$. If the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ are in the different groups, then we can add $\alpha_{v}+1$ zeros to $u_{1}$ from every second block from each group and proceed with the proof as in Case 3.2.2. Assume that blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ are in the same group, and the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ have indices $i$ and $i+hr_{2}$, respectively.
Since $\{r_{1},s_{1}\}\neq\{r_{2},s_{2}\}$, $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ are not consecutive blocks in their group. That is, $h\neq 1$ and $h\neq t-1$. In particular, $t\geqslant 4$. From each group which does not contain the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$, we can add $0^{x}$ to $u_{1}$ from every second block (from blocks with indices $i,i+2r_{2},i+4r_{2},\ldots$). If $h$ is even, then for the group containing the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ we can add $0^{x}$ to $u_{1}$ from blocks with indices $i+r_{2},i+3r_{2},i+5r_{2},\ldots$. If $h$ is odd, then for the group containing the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ we can add $0^{x}$ to $u_{1}$ from blocks with indices $i+r_{2},i+3r_{2},\ldots,i+(h-2)r_{2},i+(h-1)r_{2},i+(h+1)r_{2},i+(h+3)r_{2},i+(h+5)r_{2},\ldots$. It is not difficult to verify that for any $t\geqslant 4$ we add at most $\frac{2}{3}(a+2)$ blocks $0^{x}$ to $u_{1}$. We let $u_{1}^{\prime}$ denote the obtained subword of $u$. The proof in this case can be completed similarly to Case 3.2.2.
Case 3.2.4. $\alpha_{u}=\alpha_{v}>\beta_{u}=\beta_{v}$.